OpenStudy (anonymous):

y-c/y+c + y+c/y-c = 34/15

OpenStudy (mertsj):

What are you trying to do?

OpenStudy (anonymous):

\[\frac{ y-c }{ y+c }+\frac{ y+c }{ y-c }=\frac{ (y-c)^2+(y+c)^2 }{ y^2-c^2 }\]